With the following information make a calculation of the direct cost, direct labor and determine the base cost of raw material and the labor base cost according to the following lists.
The company KIA motors SA de CV presents the following materials:
1. #12 steel sheets, $ 1500 each.
2. High quality industrial iron, $ 45,000 per ton
3. Industrial rubber # 23 at a cost of $ 13,500 per ton
4. Electronic sensors package for a unit at a cost of $ 4,500.
5. Set of lights for one unit $ 12,500
6. Engine of 5 horses of 16 valves, 33,500 per unit
7. Game of internal accessories of the vehicle, $ 1,200 per unit package
8. Set of screws for a total assembly, $ 9,800 per game
9. Industrial nylon for protection of the units, $ 4,500 per ton
10. Transfer of finished units to the warehouse, $ 250 per unit

Labour expenses
1. 10 specialized technicians, each one receives a salary of $ 4,500
2.15 auxiliary technicians, each one receives $ 750 per day.
3. A plant supervisor with a monthly salary of $ 35,000
4. 3 auxiliary supervisors with a monthly salary of $ 15,000
5. The technicians generate a social security tax of 26% monthly on salaries.
6. A 5% monthly tax is paid to Health insurance.
7. A monthly payroll tax is paid on the technicians.

8. Finally it takes some unforeseen extra hours for auxiliary technicians of 9 hours per week.

1.Determine the base cost of the car per unit if for the case of the materials 10 cars are produced with the materials of point 2, 3 and 9 and for the case of point 1,4 steel sheets are used, in all the others a set of units.
2. Determine the base cost of labor and consider that for the construction of a car they do it in 3 hours a day.
3. Determine the total production cost if 300 cars are produced per month.
Answer #1

Base cost of Material

1. #12 Steel Sheet x 4 sheets= $ 1500 x 4= $6000
2. High Quality Industrial Iron / 10cars = $45000/10= $4500
3. Industrial rubber #23 / 10 cars = $13500/10= $1350
4. Electronic Sensor Package per unit = $4500
5. Set of lights per unit = $12,500
6.Engine of 5 horses per unit = $33500
7. Game of Internal Accessories per unit = $1200
8. Set of screws per game = $9800
9. Industrial Nylon /10 cars= $4500/10= $450
10.Transfer cost per unit = $250

Total Material Cost per car = $74050